OpenStudy (anonymous):

Solve for x and y... x + y = 3 3x - 2y = 6

OpenStudy (anonymous):

It seems y = 3/5 and x = 4/3, but I really need them to equal whole numbers...

OpenStudy (anonymous):

Am I doing something wrong with my calculations??

OpenStudy (anonymous):